A baseball is thrown with a velocity of 31.1 𝑚/𝑠 of 𝜃 = 33.4° above horizontal. What is the horizontal component of the ball’s velocity at the highest point of the ball’s trajectory?  


1
Expert's answer
2021-04-07T10:49:26-0400

To be given in question

"u" =31.1meter/sec


"\theta=33.4°"

To be asked in question

Horizontal components of highest point


"u(x)=?"

We know that

Horizontal components of velocity

"u(x)=ucos\theta"

Put value

=31.1"\times" "cos" "(33.4\u00b0)"

"u(x)" =25.96meter/sec